Two bike riders left each other and started to ride in opposite directions. Two hours later they were 54 miles apart. If one of them averaged twice the average rate of the other what was the rate of each?
Answer by nerdybill(7384)

Two bike riders left each other and started to ride in opposite directions. Two hours later they were 54 miles apart. If one of them averaged twice the average rate of the other what was the rate of each?
.
You will need to apply the "distance formula":
d = rt
where
d is distance
r is rate or speed
t is time
.
Let x = rate of one rider
then
2x = rate of other rider
.
"distance of one rider" + "distance of other rider" = "total distance"
2x + 2(2x) = 54
2x + 4x = 54
6x = 54
x = 9 mph
.
other rider:
2x = 2(9) = 18 mph
.
solution: 9 mph and 18 mph